OpenStudy (anonymous):

A building that is 115 feet tall casts a shadow that is 190 feet long. Determine the angle at which the rays of the sun hit the ground to the nearest degree.

OpenStudy (hba):

\[\Huge{\bf{WELCOME\ TO\ OS}}\]

OpenStudy (hba):



OpenStudy (ujjwal):

\[\tan \theta=\frac{AB}{BC}\]

OpenStudy (hba):

It is given that |BC|=115 feet and |AB|=190 feet We have to find \[\theta=?\]

OpenStudy (hba):

Use the formula that @ujjwal Has provided and solve for the Angle :)

OpenStudy (hba):

@Cortney95 Do you understand,if you are not clear about anything please ask ?

OpenStudy (ujjwal):

asker is absent @hba

OpenStudy (hba):

Then leave it @ujjwal Until @Cortney95 returns.

OpenStudy (anonymous):

I understand thank you so much :)

OpenStudy (hba):

Your'e welcome.

OpenStudy (anonymous):

The answer choices are A) 31° B) 37° C) 52° D) 53° and The answer I am getting is a decimal.

OpenStudy (hba):

What do you get ?

OpenStudy (hba):

@Cortney95 ?

OpenStudy (hba):

@Cortney95 Please interact ......

OpenStudy (hba):

Please round of to the nearest,if your answer is close to your options :)

OpenStudy (anonymous):

Sorry had to help family with something. My decimal was1.652173913

OpenStudy (hba):

No problem @Cortney95 ,It is better you show me your working so I can indicate where you go wrong.

OpenStudy (hba):

@Cortney95 You there ?

OpenStudy (anonymous):

I divided 190/150 and got that as a decimal since you said earlier that I should divide ab/bc and Ab=190 and Bc=115

OpenStudy (hba):

Look \[\tan \theta=\frac{ AB }{ BC }\] and \[\theta=\tan^{-1} \ (\frac{ AB }{ BC })\]

OpenStudy (anonymous):

Ohh

OpenStudy (hba):

@Cortney95 I hope now you get it :)

OpenStudy (anonymous):

I do thx

OpenStudy (hba):

Welcome,What answer did you get ?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Could you help Me with another problem?

OpenStudy (hba):

Sure please post it in another tab and tag me there :)

OpenStudy (anonymous):

Im new to all of this so I will try

OpenStudy (hba):

To do so you will have to close this question first :)

OpenStudy (anonymous):

Ok

OpenStudy (hba):

I will be back in 10 min if you post your question before that there are a lot of people here to help you.Moreover,You can help others like i helped you. Read the CoC for the rules- http://openstudy.com/code-of-conduct If you are un sure about anything ask the people in messages,you can also go the Openstudy feedback and ask question on how OS works- http://openstudy.com/study#/groups/OpenStudy%20Feedback

OpenStudy (anonymous):

Ok